OpenStudy (anonymous):

| x-3|+| x-4 |given that x>4 rewrite without using absoulute notation

OpenStudy (anonymous):

thts just x-3 + x -4 both parts are positive simplifies to 2x - 7

OpenStudy (anonymous):

what do I do with the x-4

OpenStudy (anonymous):

if x>4 , then (x-4) is positive so it's absolute value is the same . |x-4| = (x-4) Also, if x>4 , then (x-3) is positive so |x-3| = (x-3) So, tosummarize ; when x>4: |x-3|+|x-4| = (x-3)+(x-4)= x-3+x-4 = 2x-7 i.e.: |x-3|+|x-4| = 2x-7